On a History test, where the student grades are distributed normally, the average grade was 78 and the standard deviation was 10. Determine the numeric score of two students whose standard scores (a.k.a. z-scores) were -0.6 and 1.2, respectively.
Answer by stanbon(75887)

On a History test, where the student grades are distributed normally, the average grade was 78 and the standard deviation was 10. Determine the numeric score of two students whose standard scores (a.k.a. z-scores) were -0.6 and 1.2, respectively.
--------------
Since z = (x-u)/s, x = zs + u
-----
Your Problem:
x(-0.6) = -0.6*10+78 = 72
--------------
x(1.2) = 1.2*10 + 78 = 90
==============================
Cheers,
Stan H.